Jen is 15 years younger than Andre. The sum of their ages is 81. What's andres age in years?
Answer by macston(5194)

J=Jen's age=A-15; A=Andre's age
J+A=81 Substitute for J
(A-15)+A=81`
2A-15=81 Add 15 to each side
2A=96 divide each side by 2
A=48 ANSWER Andre is 48 years old
CHECK
Since Andre is 48, Jen is (48-15)=33, and 48+33=81, the given sum of their ages.